OpenStudy (anonymous):

Solve 9 - (3 - x) = 14 A. x = -8 B. x = 8 C. x - -5 D. x = 5

sam (.sam.):

9-(3-x)=14 9-3+x=14 6+x=14 x=8

OpenStudy (anonymous):

9-3+x=14 6+x=14 x=14-6 x=8 (B)

OpenStudy (anonymous):

9 - (3 - x) = 14 Or 9 - 3 + x = 14 Or 6 + x = 14 Or x = 14 - 6 Or x = 8 [B]